OpenStudy (anonymous):

When finding the mass of an equation such as Pb(NO3)2-- where is the 2 applied to the equation?

OpenStudy (lena772):

I'm assuming you mean the molar mass. Pb - 207.2g* 2 is applied to both elements in the bracket. you multiply it. N*2= 2N= 2*14= 28g* O3*2=O6=6*16=96g* add the 3 of those values (the ones with the asterisks) together and you'll have your molar mass for lead nitrate which is Pb(NO3)2

OpenStudy (lena772):

I got the molar masses of each element from the periodic table btw

OpenStudy (anonymous):

Thank you! If the question says to determine the formula mass, is molar mass what is implied?

OpenStudy (lena772):

molar mass and formula mass are the same in this case

OpenStudy (anonymous):

okay, thank you for your help!

OpenStudy (lena772):

no problem. don't hesitate to tag me & good luck!

OpenStudy (anonymous):

I want to run the whole question by you because I feel like my answer came out to be too large. Determine the formula mass of Pb(NO3)2 in g/mol and convert 0.250g Pb(N)3)2 to moles Pb(NO3)2. I did the first part of the question as you said and got an answer of 331g/mol. Then for the second part I ended up with 1324 moles. I'm new to chemistry but to me that seems like a high number of moles? I set it up as: 0.250g/1mol=331g/mol/Xmol 0.250gX=331g/mol 0.250gX/0.250gX 331g/mol/0.250gX X=1324 moles

OpenStudy (lena772):

0.250g*1mol/331g remember what you want over what you got. we want moles, we have grams. grams cancel out and we calculate. that is your answer.